OpenStudy (anonymous):

what is X^2-3x=28 my answer is (x-7)(x-4) am I correct

OpenStudy (anonymous):

No, I don't think so. If you FOIL, you can check your answer.

OpenStudy (anonymous):

what is foil :)

OpenStudy (anonymous):

First you want to have all terms on the same side of the = sign, so -28 from both sides.

OpenStudy (anonymous):

We will get to FOIL in a minute. I'll walk you through the steps to answer this question.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

\(X^2-3x=28\) First, you want the 28 on the left side of the equal sign.

OpenStudy (anonymous):

following keep going :)

OpenStudy (anonymous):

\(x^2-3x-28=0\) To factor, start with the \(x^2\) term. This one is easy since the coefficient is 1. (x )(x )

OpenStudy (anonymous):

Now you need to find 2 numbers that multiply to be -28 and add to be -3

OpenStudy (anonymous):

What do you think they are?

OpenStudy (anonymous):

-7 and +4

OpenStudy (anonymous):

Yes, now you can write your factors....

OpenStudy (anonymous):

x-7 x +4

OpenStudy (anonymous):

x-7 x+4

OpenStudy (anonymous):

\((x+4)(x-7)\) And to check your work - use FOIL First = \(x \times x = x^2\) Outside = \(-7 \times x = -7x\) Inside = \(4 \times x = 4x\) Last = \(-7 \times 4 = -28\)

OpenStudy (anonymous):

Put it all together and you should get back to your original equation.

OpenStudy (anonymous):

neat concept awesome you get a medal for that :) thnx bunches

OpenStudy (anonymous):

Your welcome.